OpenStudy (anonymous):

@Demonx341 @Angel_Bear19 The conservation club has 32 members. There are 18 girls in the club. What is the ratio of girls to boys in simplest form? A. 7/16 B. 9/16 C. 7/9 D. 9/7

OpenStudy (anonymous):

hi chocobo

OpenStudy (darkbluechocobo):

32-18=14 so 14 guys

OpenStudy (darkbluechocobo):

and hello there

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the fraction 14/18 is not reduced to lowest terms. We can reduce this fraction to lowest terms by dividing both the numerator and denominator by 2. Why divide by 2? 2 is the Greatest Common Divisor (GCD) or Greatest Common Factor (GCF) of the numbers 14 and 18. So, this fraction reduced to lowest terms is 7/9 so the answer is C.

OpenStudy (anonymous):

oh ok

OpenStudy (darkbluechocobo):

18/32-14/18= 4/18 2/9

OpenStudy (darkbluechocobo):

buht ayeee im wrong

OpenStudy (darkbluechocobo):

can confirm i suck a ratios lel

OpenStudy (anonymous):

oh lol

OpenStudy (darkbluechocobo):

but demon luckily knows